OpenStudy (anonymous):

What is the length of a diagonal brace that could be used for a wall 9 feet high and 12 feet long? 15 ft 14 ft 13 ft 12 ft

OpenStudy (mertsj):

Try using the pythagorean theorem

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

so 9^2+12^2=c^2

OpenStudy (anonymous):

you can do the rest

OpenStudy (mertsj):

That's it

OpenStudy (anonymous):

thaks

OpenStudy (anonymous):

no problem